OpenStudy (anonymous):

\[ \frac{e^{2y}-1}{e^{2y}+1} = x \] How can we solve for e^y ?

OpenStudy (anonymous):

First Cross Multiply

OpenStudy (anonymous):

e^2y - 1 = x e^2y + x e^2y - x e^2y = x + 1 e^2y ( 1-x) = x + 1 e^2y = (x+1) / (1-x) e^2y = (e^y)^2 Take sqrt on both sides and u have e^y..:)

OpenStudy (anonymous):

ie e^y = sqrt [ (x+1) / (1-x) ]